OpenStudy (anonymous):

3.5x + 5 - 1.5x = 8 ???

OpenStudy (anonymous):

3.5x-1.5x=8-5 2x/2=3/2 x=3/2

OpenStudy (merchandize):

2.5x=8-5 2.5x=3 x=3/2.5 x=1.2

OpenStudy (anonymous):

wut.